Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.25 coming 3 years from today. The dividend should grow rapidly - at a rate of 50% per year - during Years 4 and 5, but after Year 5, growth should be a constant 4% per year. If the required return on Computech is 16%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Dividend year 3 =1.25
Dividend year 4 =D3*(1+50%)=1.25*1.5
Dividend Year 5 =D4*(1+50%)=1.25*1.5^2
Dividend Year 6=D5*(1+4%)=1.25*1.5^2*1.04
Terminal Year 5=1.25*1.5^2*1.04/(16%-4%) =24.3750
Value of the stock today =Dividend 3/(1+r)^3+Dividend Year 4/(1+r)^4+Dividend 5/(1+r)^5+Terminal Value/(1+r)^5
=1.25/1.16^3+1.25*1.5/1.16^4+1.25*1.5^2/1.16^5+24.3750/1.16^5 =14.78